ganeshie8 (ganeshie8):

\(d = ax + by\) \(d\) is the smallest whole number; does this mean \(d = \gcd(a, b) = \gcd(x, y)\) ?

ganeshie8 (ganeshie8):

i was looking at gcd(a, b) proof and got stuck on this doubt...

OpenStudy (ikram002p):

means there is unique d s.t d=ax+by

ganeshie8 (ganeshie8):

yes is \(d\) also the gcd of \(x\) and \(y\) ?

OpenStudy (ikram002p):

nope its not

ganeshie8 (ganeshie8):

I thought so, but if d is smallest whole number then it has to be the gcd right ?

OpenStudy (ikram002p):

lets prove it ^

ganeshie8 (ganeshie8):

okay..

myininaya (myininaya):

Counterexample gcd(14,10)=2 3(10)-14(2)=2 gcd(3,2)=1 not 2

OpenStudy (ikram002p):

it has to be GCD of a,b since function u have is s:={ au+bv | s.t au+bv \(\in N\)}

OpenStudy (ikram002p):

exactly :) @myininaya

ganeshie8 (ganeshie8):

oh we cannot swap (a, b) and (x, y) is it, its still hard to see..

OpenStudy (ikram002p):

well if u swap x,y then u need to swap the wole proof as well and they would be other d1=gcd(x,y) s.t \(d_1\) = mx+ny

ganeshie8 (ganeshie8):

I see... the counter example clears up everything xD thank you both for quick help !

OpenStudy (ikram002p):

np :)

myininaya (myininaya):

also -11(10)+8(14)=2 but gcd(10,14)=2 and gcd(-11,8)=gcd(11,8)=1 i wonder if we can find two numbers a,b so that 10a+14b=2 and gcd(a,b) is not equal to 1 so far I can only find gcd(a,b) is 1

ganeshie8 (ganeshie8):

d = ax+by can have multiple representations is it ?

ganeshie8 (ganeshie8):

-11(10)+8(14)=2 i feel (-11, 8) is the unique solution for 10x + 14y = 2 just a guess, i may be very wrong...

ganeshie8 (ganeshie8):

my bad ! 10x + 14y = 2 is a diaphontine equation, it can have infinite solutions

myininaya (myininaya):

yeah one solution i wrote as (3,-2) earlier

OpenStudy (ikram002p):

its a condition that gcd(x,y)=1 disrigarding d value

OpenStudy (ikram002p):

i only need to make sure of that mm not sure yet

ganeshie8 (ganeshie8):

since (3, -2) is one solution, all other solutions can be given by : x = 3 + 7t y = -2 - 5t

ganeshie8 (ganeshie8):

that means gcd(3+7t, -2-5t) is always 1 for all values of t @ikram002p ?

OpenStudy (ikram002p):

i think so

myininaya (myininaya):

right thanks for some reason i was having a brain fart on how to find that so gcd(x,y)=gcd(3+7t,-2-5t) This might equal 1 I'm just wondering if for all x,y will the gcd(x,y)=1 where the 10x+14y=2. so gcd(3+7t,-2-5t) =gcd(3+7t,1+2t) =gcd(-t,1+2t) =gcd(t+1,1+2t) =gcd(t+1,t) =1 for all t

myininaya (myininaya):

you know since no two consecutive integers will have a common factor

OpenStudy (ikram002p):

nice !

OpenStudy (ikram002p):

@ganeshie8 you rivewing Burton book ? there is a collary after thm 2.3 check it

myininaya (myininaya):

I think I made a mistake up there gcd(3+7t,-2-5t) =gcd(3+7t,1+2t) =gcd(t,1+2t) <--not -t but putting -t or t I don't think should actually matter =gcd(t,1) =1 for all t

ganeshie8 (ganeshie8):

wow! this looks interesting xD you're talking about this ikram ? http://prntscr.com/40fm9c

OpenStudy (ikram002p):

yep .. that make me really think for any x,y gcd(x,y)=1 ax+by=d then a(mg)+b(mh)=d ag+bh=d/m then d wont be the least whol number according to WOP

ganeshie8 (ganeshie8):

another way to see gcd(x, y) has to be 1 is : ax + by = gcd(a, b) dividing both sides by gcd(a, b) gives : px + qy = 1 where p, q are coprime, So gcd(x, y) = 1

myininaya (myininaya):

http://www.saylor.org/site/wp-content/uploads/2013/05/An-Introductory-in-Elementary-Number-Theory.pdf like theorem 8 on page 21 of this pdf

myininaya (myininaya):

that is what i was using above

myininaya (myininaya):

ok yeah @ganeshie8

myininaya (myininaya):

that is probably easier

OpenStudy (ikram002p):

cool

OpenStudy (ikram002p):

@ganeshie8 mmm why x,y are coprime when ax+by =1 ?

ganeshie8 (ganeshie8):

Recall : a = bc => b|a and c|a

ganeshie8 (ganeshie8):

px+qy = 1 if x,y are NOT coprime, then you can factor out gcd(x,y) then this gcd has to divide 1 which is impossible.

OpenStudy (ikram002p):

aha nice :)